OpenStudy (anonymous):

6x^2+15y^2=16 I don't know what to do with the y^2

OpenStudy (jdoe0001):

anything wrong with it? looks like a well-behaving "y", so is a nice "y", is really a lovely one

OpenStudy (anonymous):

y^2=-6/15x^2+16/15 I don't know what to do.

OpenStudy (jdoe0001):

\(\bf 6x^2+15y^2=16\implies y^2=-\cfrac{6x^2}{15}+\cfrac{16}{5} \\ \quad \\ taking\implies \sqrt{\qquad }\qquad thus \\ \quad \\ \sqrt{y^2}=\sqrt{-\cfrac{6x^2}{15}+\cfrac{16}{5} }\implies y=\sqrt{-\cfrac{6x^2}{15}+\cfrac{16}{5} }\)